OpenStudy (anonymous):

7th grade math help plzzz

OpenStudy (anonymous):

The formula n = rt gives the number n of boxes packed in time t at rate r. A machine packs boxes at a constant rate of 11.5 boxes per minute. How many boxes will the machine pack in 6 min?

OpenStudy (anonymous):

@aum

OpenStudy (cj49):

11.5x6=?

OpenStudy (anonymous):

69???

OpenStudy (anonymous):

could u help me with 4 more please:)

OpenStudy (cj49):

sure

OpenStudy (anonymous):

Solve the word problem. The base of a triangle with height h and area A is given by the formula b = 24/h . What is the base of a triangle with area 17 cm2 and height 5 cm?

OpenStudy (anonymous):

A. 3.4 cm B. 3.8 cm C. 6.8 cm D. 11 cm

OpenStudy (cj49):

6.8

OpenStudy (anonymous):

Tony wants to calculate his average quiz score. The formula A = T/n gives Tony’s average score A for a total of T points on n quizzes. Tony had 47 points on 5 quizzes. What is Tony’s average score?

OpenStudy (cj49):

47/5=?

OpenStudy (anonymous):

Wrong section to post this in.

OpenStudy (anonymous):

9.4????

myininaya (myininaya):

Hey @cjackson2002 you would get more help if you post this in the math section I bet. Though I know we have had problems or I have seen problems where somehow questions that are posted in the right section get posted elsewhere according to other people's OpenStudy view.

OpenStudy (anonymous):

Indeed.

myininaya (myininaya):

I remember people claiming there was a question in math while it was in feedback.

OpenStudy (cj49):

yes its 9.4

OpenStudy (anonymous):

Solve the problem. The formula d = rt gives the distance traveled in time t at rate r. If a bicyclist rides for 2.3 hours and travels 41 miles, what was her average speed? A. 9.43 miles per hour B. 5.6 miles per hour C. 17.83 miles per hour D. 38.7 miles per hour

OpenStudy (cj49):

ill suggest u to post this questions in the maths section as mentioned by @myininaya and @Basher

OpenStudy (anonymous):

kk